OpenStudy (anonymous):

Which of the following options results in a graph that shows exponential growth? f(x) = 0.4(0.2)x f(x) = 4(0.98)x f(x) = 0.7(5)x f(x) = 5(4)−x

OpenStudy (anonymous):

Look at the number inside the parenthesis; if it is greater than 1, then that's an exponential growth.

OpenStudy (anonymous):

Examples: \(\huge f(x)=100(2)^x\) \(\huge f(x)=0.5(9)^x\) The numbers in the parentheses are 2 and 9. They are greater than 1, so they both show an exponential growth.

OpenStudy (anonymous):

Did you understand it?